OpenStudy (anonymous):

2/3 of what number is 6?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

i know that i need to know how to do the work to the problem.

OpenStudy (anonymous):

2/3 multiplied by x = 6 So x = 6 * 3/2 = 9